# ANSR-DT: An Adaptive Neuro-Symbolic Learning and Reasoning Framework for Digital Twins thanks: This research was partially supported by the U.S. National Science Foundation through Grant Nos. 2317117 and 2309760. (Corresponding author: Safayat Bin Hakim.)

### v2

@@ Abstract
Index Terms:
Digital Twin, Neuro-Symbolic AI, Adaptive Intelligence, Human-Machine Collaboration, Reinforcement Learning, CNN-LSTM
@@ I Introduction
ANSR-DT combines neural networks, symbolic reasoning, and reinforcement learning to create adaptive, interpretable digital twins for industrial use, achieving high accuracy and stable rule extraction with a refined framework tested on a synthetic dataset.
@@
Digital Twins (DTs), which are real-time digital replicas of physical systems, play an important role in bridging the gap between the physical and digital worlds [2, 3]. They help improve monitoring, decision-making, and optimization in complex scenarios such as smart factories, autonomous systems, and various adaptable environments [4]. However, traditional Digital Twins (DTs) often struggle with adaptability and interpretability, limiting trust [5]. While Dynamic Digital Twins (DDTs) and Dynamic Data-Driven Application Systems (DDDAS) attempt to address these challenges through the integration of real-time data [6], they still struggle to seamlessly incorporate human input and ensure the interpretability of complex decision processes [7].
To address these challenges, we propose an Adaptive Neuro-Symbolic Learning Framework for digital twin technology called ANSR-DT. This framework uses Proximal Policy Optimization (PPO) algorithm [8] in collaboration with the CNN-LSTM and attention technique [9] to ensure logical clarity of symbolic reasoning [10], aiming to create an adaptive system that maintains interpretability while addressing user needs.
The key contributions of ANSR-DT include:
  • We propose a hybrid neuro-symbolic architecture for digital twin technology that integrates deep learning and symbolic reasoning to improve decision-making processes in operational environments, addressing the interpretability gap in current DT implementations [11].
  • The PPO algorithm in collaboration with the CNN-LSTM, within the proposed ANSR-DT framework, enables continuous learning that ensures real-time adaptation to user preferences and environmental changes [12].
  • The statistical results from the considered scenarios demonstrate strong performance in terms of comparative metrics, with peak validation accuracy of 99.5% for dynamic event detection. In addition, symbolic reasoning used in this work is a step forward in developing real-world algorithms that yield logically reasoned results for better operation and performance.
  • An open-source implementation that facilitates reproducibility and invites future research in adaptive digital twins.
@@
The key technical challenge addressed by ANSR-DT is the integration of real-time adaptability and interpretability in digital twins, which traditional models fail to achieve due to their reliance on static rule sets and lack of continuous learning mechanisms [13]. Current approaches suffer from three critical limitations: (1) inability to adapt to real-time human input with response latencies under 1 second, (2) black-box decision processes that lack transparency for human operators, and (3) static symbolic representations that cannot evolve with changing operational conditions. ANSR-DT overcomes these limitations through a novel architecture that combines neural networks for dynamic event detection, symbolic reasoning for interpretable decision-making, and reinforcement learning for continuous adaptation.
@@ II Background and Related Work
Unlike traditional digital twins that use static models [14], ANSR-DT integrates neural networks and symbolic reasoning for adaptability and interpretability, building on neuro-symbolic advances [15, 16, 17] and reinforcement learning’s dynamic capabilities [18, 19], tailored to industrial needs with a streamlined dataset and extended training.
Traditional DT implementations face three common limitations: (1) reliance on predefined rules that cannot evolve with changing operational conditions, (2) limited interpretability of decision processes, and (3) insufficient integration of human feedback [20]. While they improve upon static digital models, they still fail to address the core challenges of adaptive intelligence in complex industrial settings.
The limitations of current approaches can be summarized as:
  1. 1.
    Lack of Continuous Learning: Existing approaches typically employ fixed rule sets or models that cannot adapt to changing conditions without manual intervention [21].
  2. 2.
    Poor Interpretability: Many current systems operate as black boxes, making their decisions difficult to understand or trust by human operators [22].
  3. 3.
    Limited Human Integration: Most frameworks do not effectively incorporate human feedback for continuous improvement.
  4. 4.
    Insufficient Adaptability: Current digital twins struggle to respond quickly to dynamic changes in the physical environment [5].
  5. 5.
    Isolated Techniques: Approaches tend to focus on either neural learning or symbolic reasoning, but rarely integrate them effectively [23].
To address these limitations, we propose ANSR-DT, a framework that combines adaptive neuro-symbolic reasoning and real-time reinforcement learning to enable dynamic and interpretable decision-making in digital twin applications.
@@ III-A ANSR-DT Architecture Overview
The proposed ANSR-DT framework integrates neural networks, symbolic reasoning, and reinforcement learning to create an adaptive, interpretable digital twin system. As illustrated in Fig. 1, ANSR-DT employs a three-layer architecture consisting of: (1) Physical Layer for sensor integration, (2) Processing Layer for neural-symbolic reasoning, and (3) Adaptation Layer for continuous learning and rule updating. This architecture facilitates seamless integration between physical systems and their digital representations while maintaining interpretability and adaptability.
Each layer serves a specific function while interacting with others through well-defined interfaces. The Physical Layer collects and preprocesses sensor data; the Processing Layer translates these signals into meaningful patterns and symbolic rules; and the Adaptation Layer continuously refines the system’s decision-making policies based on performance feedback and changing conditions. Fig. 2 illustrates the operational sequence and information flow between these layers.
@@ III-B Physical Layer: Sensor Integration
The Physical Layer forms the foundation of ANSR-DT, connecting the digital twin to its physical counterpart through multiple sensor types:
  • Temperature Sensors monitor thermal variations in industrial equipment, detecting potential overheating conditions or thermal deviations
  • Vibration Sensors measure mechanical oscillations, providing early warnings of component wear or imbalance
  • Pressure Sensors track fluid dynamics within systems, identifying potential leaks or flow restrictions
Raw sensor data undergoes systematic preprocessing before being passed to the Processing Layer. This includes noise filtering, normalization, and temporal alignment to ensure consistent interpretation across different sensor types and sampling rates [24].
@@ III-C Processing Layer: Neural-Symbolic Reasoning
The Processing Layer constitutes the cognitive core of ANSR-DT, combining deep learning for time-series pattern extraction with symbolic reasoning for interpretable decision-making.
@@ III-C1 Neural Component: CNN-LSTM Architecture
The CNN-LSTM comprises two Conv1D layers with 64 and 128 filters (kernel size 3), followed by batch normalization and max-pooling, then two LSTM layers with 100 and 50 units, respectively, and an attention mechanism to prioritize critical time steps. This configuration totals 189,481 parameters, balancing complexity and efficiency for the 5,000-sample dataset.
This neural component transforms raw sensor data into feature vectors representing detected patterns, key events, or state changes. These vectors then serve as inputs to the symbolic reasoning engine.
@@ III-C2 Symbolic Component: Rule-Based Reasoning
The symbolic reasoning engine translates neural outputs into interpretable logical rules and facts [25]. This component maintains a knowledge base that represents system states, operational constraints, and decision logic in first-order predicate logic format.
The symbolic component provides several key advantages:
  • Human-readable representation of system knowledge
  • Explicit encoding of domain expertise and operational constraints
  • Traceable decision paths from observations to actions
  • Ability to incorporate human feedback and preferences
The integration between neural and symbolic components occurs through a rule extraction process governed by Equation 1:
$R_{\text{new}}=\{r|\text{conf}(r)>\tau,r\in\mathcal{F}(M_{\theta},\mathcal{D})\}$ (1)
In this equation, new rules $R_{\text{new}}$ are selected if they meet two criteria: (1) they are derived from neural model outputs using feature extraction function $\mathcal{F}$ with model parameters $M_{\theta}$ and data $\mathcal{D}$, and (2) their confidence score $\text{conf}(r)$ exceeds threshold $\tau$. This ensures that only reliable and relevant rules are incorporated into the knowledge base.
In the implementation, the SymbolicReasoner component performs this integration through the extract_rules_from_neural_model function, which analyzes patterns detected by the neural network and converts them into symbolic rules. The update_rules method then incorporates these rules into the Prolog rule base for logical inference [26].
@@ III-D Adaptation Layer: Reinforcement Learning
The Adaptation Layer enables ANSR-DT to continuously improve its performance through reinforcement learning, specifically using the Proximal Policy Optimization (PPO) algorithm.
@@ III-D1 PPO Algorithm Implementation
The PPO algorithm optimizes the policy for decision-making while ensuring stable updates through a clipped objective function, as shown in Equation 2:
@@
Where $r_{t}(\theta)=\frac{\pi_{\theta}(a_{t}\mid s_{t})}{\pi_{\theta_{\text{old}}}(a_{t}\mid s_{t})}$ represents the probability ratio of selecting action $a_{t}$ in state $s_{t}$ under the current policy compared to the previous policy. The Advantage function $A_{t}=Q(s_{t},a_{t})-V(s_{t})$ measures the relative value of taking specific actions. The entropy term $H[\pi_{\theta}]$ promotes exploration, with its impact controlled by coefficient $\beta=0.01$.
In our implementation, we set $\beta=0.01$ to balance exploration with exploitation, and use a clipping range $\epsilon=0.2$ to limit policy updates. These parameters were selected to ensure stable learning while maintaining adaptability in dynamic industrial environments.
Extending PPO training from 5 iterations (10,240 timesteps) to 98 iterations (200,704 timesteps) raised the explained variance from 0.447 to 0.547, improving long-term prediction accuracy. Using consistent hyperparameters (learning rate 1e-3, clip_range 0.2), this extended training enhanced the agent’s ability to anticipate and adapt to dynamic industrial conditions, as evidenced by refined actions in inference scenarios.
We optimize the PPO implementation with the following parameter settings:
  • Batch size: 64 samples for stable gradient estimation
  • Learning rate: $10^{-3}$ for efficient policy updates
  • Training epochs: 10 per policy update
  • Discount factor: $\gamma=0.99$ for balancing immediate and future rewards
  • Clipping parameter: $\epsilon=0.2$ to limit update size
  • GAE parameter: $\lambda=0.95$ for advantage estimation
@@ III-D2 Adaptive Exploration Strategy
To balance exploration and exploitation, we implement an adaptive context-aware exploration approach governed by Equation 3:
@@
This strategy dynamically adjusts the system’s behavior based on the current operational context. When the system detects a point of interest through sensor data analysis and symbolic reasoning, it switches to exploratory mode by setting the deterministic flag to false, encouraging the discovery of new optimal actions. During normal operation, the system defaults to exploitative mode, prioritizing actions with known high rewards [19].
@@ III-D3 Dynamic Rule Updating
The Adaptation Layer continuously refines the symbolic knowledge base by incorporating new rules derived from reinforcement learning outcomes. This dynamic rule updating process ensures that the symbolic reasoning component evolves alongside changing operational conditions and user preferences.
@@ III-E Framework Integration and Adaptive Operation
ANSR-DT implements tight integration across its three layers to enable seamless information flow and adaptive decision-making. Fig. 2 illustrates the operational sequence and component interactions during system operation.
III-E1 Physical-Digital Integration
The framework establishes bidirectional communication between physical and digital components through:
  • Data Preprocessing Pipeline: Sensor data undergoes systematic validation and normalization before feeding into the processing layer, with specific steps:
    • Standardization of sensor readings using rolling statistics
    • Temporal alignment of multi-sensor data streams
    • Identification of deviations in raw sensor readings
  • Feedback Control Loop: The processing layer provides real-time feedback for sensor network optimization:
    • Dynamic adjustment of sampling rates based on system state
    • Automated sensor calibration based on performance metrics
    • Fault detection and sensor reliability assessment
III-E2 Continuous Adaptation Mechanism
The system maintains continuous adaptation through an integrated feedback loop encompassing two key processes:
  • State Assessment: Continuous monitoring of:
    • System performance metrics
    • Efficiency indices
    • Rule activation patterns and effectiveness
    • Policy performance across operational modes
  • Adaptive Response: Coordinated system adjustments including:
    • Policy refinement through PPO-based learning
    • Knowledge base evolution via dynamic rule extraction
    • Sensor network configuration optimization
This integration ensures that ANSR-DT maintains optimal performance while adapting to changing operational conditions, with each layer contributing to the system’s overall adaptability and robustness [27].

### v3

@@ I Introduction
Digital Twins (DTs), real-time digital replicas of physical systems, bridge the gap between physical and digital worlds to improve monitoring, decision-making, and optimization in complex scenarios such as smart factories, autonomous systems, and adaptable environments [2, 3, 4]. Despite their promise, traditional DT implementations struggle with adaptability and interpretability, fundamentally limiting trust in safety-critical industrial settings [5]. While Dynamic Digital Twins (DDTs) and Dynamic Data-Driven Application Systems (DDDAS) attempt to address these challenges through real-time data integration [6], they remain constrained by their inability to seamlessly incorporate human input and ensure transparent decision processes [7].
@@
The core technical challenge lies in integrating real-time adaptability with interpretability—traditional models fail because they rely on static rule sets and lack continuous learning mechanisms [8]. Current approaches suffer from three critical limitations: inability to adapt to real-time human input with response latencies under 1 second, opaque decision processes that lack transparency for human operators, and static symbolic representations that cannot evolve with changing operational conditions. Existing solutions either prioritize neural learning or symbolic reasoning, but rarely integrate them effectively to achieve both high performance and interpretability [9].
To overcome these fundamental limitations, we propose ANSR-DT, an Adaptive Neuro-Symbolic Learning and Reasoning Framework for digital twin technology. ANSR-DT combines CNN-LSTM neural networks for dynamic event detection, Prolog-based symbolic reasoning for interpretable decision-making, and Proximal Policy Optimization (PPO) reinforcement learning [10] for continuous adaptation. This integration leverages attention mechanisms [11] to prioritize critical temporal patterns while maintaining logical clarity through symbolic reasoning [12], creating an adaptive system that addresses user needs without sacrificing interpretability.
The framework makes four principal contributions to adaptive digital twin technology:
  • A hybrid neuro-symbolic architecture that integrates deep learning and symbolic reasoning to improve decision-making processes in operational environments, addressing the interpretability gap in current DT implementations [13].
  • A continuous learning mechanism through PPO-enhanced CNN-LSTM integration that ensures real-time adaptation to user preferences and environmental changes [14].
  • Demonstrated performance improvements over traditional approaches, achieving 99.5% peak validation accuracy for dynamic event detection while extracting 14 stable symbolic rules that provide logically reasoned, human-interpretable explanations.
  • An open-source implementation that facilitates reproducibility and establishes a foundation for future research in adaptive, interpretable digital twins.
The remainder of this paper proceeds as follows. Section II examines related work and positions ANSR-DT within the current research landscape. Section III details the three-layer architecture comprising physical sensor integration, neuro-symbolic processing, and reinforcement learning adaptation. Section IV describes the technical realization including synthetic data generation and component integration. Section V presents comprehensive experimental evaluation demonstrating performance improvements and adaptability enhancements. Section VI analyzes the implications and limitations of our approach. Section VII concludes with future research directions.
@@ II Background and Related Work
Traditional digital twin implementations face persistent challenges that limit their effectiveness in dynamic industrial environments. Unlike static digital models [15], contemporary DT systems require both adaptability to changing conditions and interpretability for human operators—requirements that existing approaches struggle to satisfy simultaneously. While recent work in neuro-symbolic AI [16, 17, 18] and reinforcement learning [19, 20] offers promising directions, these advances have not been systematically integrated into digital twin architectures tailored for industrial applications.
Current DT implementations suffer from fundamental architectural limitations. Most systems rely on predefined rules that cannot evolve with changing operational conditions, requiring manual intervention whenever environmental dynamics shift beyond initial specifications [21]. This rigidity becomes problematic in industrial settings where process parameters, equipment characteristics, and operational objectives continuously evolve. Furthermore, many state-of-the-art systems operate as opaque decision-makers, making their reasoning processes difficult to understand or validate [22]. This opacity directly undermines trust, particularly when human operators must verify system recommendations before implementing critical adjustments. The limited integration of human feedback exacerbates these challenges—without effective mechanisms to incorporate operator expertise and preferences, DT systems cannot align their behavior with organizational knowledge accumulated over years of operation [23].
The adaptability deficit manifests most clearly during operational deviations. Current digital twins struggle to respond quickly to dynamic changes in the physical environment, often failing to detect subtle precursors to critical failures or recognize complex multi-sensor patterns indicating emerging problems [5]. This limitation stems partly from the isolated application of techniques—approaches tend to focus exclusively on either neural learning for pattern recognition or symbolic reasoning for rule-based inference, failing to leverage the complementary strengths of both paradigms [9]. Neural networks excel at extracting complex patterns from high-dimensional sensor data but provide little insight into their decision-making processes. Conversely, symbolic systems offer transparent reasoning but lack the flexibility to adapt to novel patterns not explicitly encoded in their rule bases.
Recent neuro-symbolic AI research demonstrates the potential of hybrid architectures that combine neural learning with symbolic reasoning. Schmidt et al. [16] systematically examined neuro-symbolic approaches for knowledge graph construction in manufacturing, identifying key integration patterns between subsymbolic and symbolic representations. Zhou et al. [17] demonstrated practical neuro-symbolic AI deployment at Bosch, highlighting the importance of data foundations and systematic integration strategies. These works establish that effective neuro-symbolic integration requires careful attention to the bidirectional flow of information between neural and symbolic components—a principle central to ANSR-DT’s architecture.
ANSR-DT builds on these foundations by integrating neural networks and symbolic reasoning specifically for digital twin adaptability and interpretability. The framework extends beyond existing neuro-symbolic approaches through its incorporation of reinforcement learning for continuous policy adaptation, enabling the system to refine both its neural pattern recognition and symbolic rule base based on observed outcomes. This three-way integration—neural learning, symbolic reasoning, and reinforcement learning—distinguishes ANSR-DT from prior work while addressing the core limitations of traditional digital twin implementations. By maintaining a streamlined dataset design and extended training protocols, the framework balances computational efficiency with the depth of adaptation required for industrial deployment.
@@ III-A ANSR-DT Architecture Overview
ANSR-DT integrates neural networks, symbolic reasoning, and reinforcement learning through a three-layer architecture designed to provide both adaptability and interpretability in digital twin applications. As illustrated in Fig. 1, the framework comprises a Physical Layer for sensor integration, a Processing Layer for neuro-symbolic reasoning, and an Adaptation Layer for continuous learning and rule updating. This architecture facilitates seamless integration between physical systems and their digital representations while maintaining transparent decision processes that human operators can understand and validate.
@@
Each layer serves a distinct function while maintaining well-defined interfaces with adjacent layers. The Physical Layer collects and preprocesses sensor data, transforming raw measurements into standardized inputs for downstream processing. The Processing Layer translates these signals into meaningful patterns through neural networks and converts detected patterns into human-readable symbolic rules through logical inference. The Adaptation Layer continuously refines the system’s decision-making policies based on performance feedback and changing operational conditions, ensuring that both neural and symbolic components evolve together. The operational sequence and information flow between these layers is detailed in Fig. 2, which shows how sensor data progresses through pattern detection, symbolic reasoning, and adaptive control decisions.
@@ III-B Physical Layer: Sensor Integration
The Physical Layer establishes the foundation for ANSR-DT by connecting the digital twin to its physical counterpart through multiple sensor modalities:
  • Temperature Sensors monitor thermal variations in industrial equipment, detecting potential overheating conditions or thermal deviations that may indicate degraded performance or impending component failures.
  • Vibration Sensors measure mechanical oscillations, providing early warnings of component wear, imbalance, or structural degradation through characteristic frequency signatures.
  • Pressure Sensors track fluid dynamics within systems, identifying potential leaks, flow restrictions, or anomalous pressure buildup that could compromise operational safety.
Raw sensor data undergoes systematic preprocessing before entering the Processing Layer. This preprocessing pipeline includes noise filtering to remove measurement artifacts, normalization to ensure consistent scaling across sensor types, and temporal alignment to synchronize multi-sensor data streams with potentially different sampling rates [24, 25]. The standardized sensor outputs then flow into the neural component for pattern extraction, establishing a consistent data foundation for downstream reasoning.
@@ III-C Processing Layer: Neural-Symbolic Reasoning
The Processing Layer constitutes the cognitive core of ANSR-DT, combining deep learning for temporal pattern extraction with symbolic reasoning for interpretable decision-making. This dual-component architecture addresses the fundamental tension between learning capacity and explainability that limits many contemporary AI systems.
@@ III-C1 Neural Component: CNN-LSTM Architecture
The neural component employs a CNN-LSTM architecture comprising two Conv1D layers with 64 and 128 filters (kernel size 3), followed by batch normalization and max-pooling to extract spatial features. Two LSTM layers with 100 and 50 units then capture temporal dependencies across sensor sequences, while an attention mechanism prioritizes critical time steps within the temporal patterns. This configuration totals 189,481 parameters, carefully balanced to provide sufficient model capacity for the 5,000-sample dataset without risking overfitting on the relatively compact training set.
The neural component transforms raw sensor data into feature vectors representing detected patterns, critical events, or state changes. These vectors capture complex multivariate relationships across temperature, vibration, and pressure measurements that would be difficult to encode through manual feature engineering. The learned representations then serve as inputs to the symbolic reasoning engine, establishing a bridge between subsymbolic pattern recognition and symbolic logical inference.
@@ III-C2 Symbolic Component: Rule-Based Reasoning
The symbolic reasoning engine translates neural outputs into interpretable logical rules and facts using a Prolog-compatible representation [26]. This component maintains a knowledge base representing system states, operational constraints, and decision logic in first-order predicate logic format. The symbolic representation provides human-readable encoding of system knowledge, explicit representation of domain expertise and operational constraints, traceable decision paths from observations to actions, and direct mechanisms to incorporate human feedback and preferences into the reasoning process.
The integration between neural and symbolic components follows the rule extraction process governed by Eq. 1. New rules $R_{\text{new}}$ are selected when they meet two criteria: derivation from neural model outputs through feature extraction function $\mathcal{F}$ operating on model parameters $M_{\theta}$ and data $\mathcal{D}$, and confidence scores $\text{conf}(r)$ exceeding threshold $\tau$. This dual-threshold approach ensures that only reliable and relevant rules enter the knowledge base.
$R_{\text{new}}=\{r\mid\text{conf}(r)>\tau,r\in\mathcal{F}(M_{\theta},\mathcal{D})\}$ (1)
In the implementation, the SymbolicReasoner component performs this integration through the extract_rules_from_neural_model function, which analyzes patterns detected by the neural network and converts them into symbolic rules. The update_rules method then incorporates these rules into the Prolog rule base for logical inference [27]. This bidirectional interface ensures that insights from neural learning inform symbolic reasoning while symbolic constraints guide neural attention toward operationally relevant patterns.
@@ III-D Adaptation Layer: Reinforcement Learning
The Adaptation Layer enables ANSR-DT to continuously improve its performance through reinforcement learning, specifically using the Proximal Policy Optimization (PPO) algorithm to refine decision-making policies while ensuring stable updates.
@@ III-D1 PPO Algorithm Implementation
The PPO algorithm optimizes the policy for decision-making through a clipped objective function that balances learning progress with stability, as formalized in Eq. 2. The probability ratio $r_{t}(\theta)=\frac{\pi_{\theta}(a_{t}\mid s_{t})}{\pi_{\theta_{\text{old}}}(a_{t}\mid s_{t})}$ compares the current policy’s action selection probability to the previous policy’s probability for action $a_{t}$ in state $s_{t}$. The Advantage function $A_{t}=Q(s_{t},a_{t})-V(s_{t})$ quantifies the relative value of taking specific actions compared to the average value of state $s_{t}$. The entropy term $H[\pi_{\theta}]$ promotes exploration of the action space, with its impact controlled by coefficient $\beta=0.01$ to balance exploration with exploitation of known good policies.
@@
Our implementation employs $\beta=0.01$ to balance exploration with exploitation and uses a clipping range $\epsilon=0.2$ to limit the magnitude of policy updates, preventing destabilizing changes. These parameters were selected to ensure stable learning while maintaining sufficient adaptability for dynamic industrial environments. Extending PPO training from 5 iterations (10,240 timesteps) to 98 iterations (200,704 timesteps) improved explained variance from 0.447 to 0.547, demonstrating enhanced long-term prediction accuracy. This extended training, conducted with consistent hyperparameters (learning rate $10^{-3}$, clip range 0.2), strengthened the agent’s ability to anticipate and adapt to dynamic industrial conditions, as evidenced by refined control actions in operational inference scenarios.
The PPO implementation employs a batch size of 64 samples for stable gradient estimation, a learning rate of $10^{-3}$ for efficient policy updates, 10 training epochs per policy update, a discount factor $\gamma=0.99$ for balancing immediate and future rewards, a clipping parameter $\epsilon=0.2$ to limit update magnitude, and a GAE parameter $\lambda=0.95$ for advantage estimation. These settings collectively ensure that policy learning proceeds steadily without catastrophic forgetting or excessive oscillation in performance.
@@ III-D2 Adaptive Exploration Strategy
ANSR-DT implements an adaptive context-aware exploration approach that dynamically adjusts behavior based on operational context, as formalized in Eq. 3. When the system detects deviations through sensor analysis and symbolic reasoning, it switches to exploratory mode by setting the deterministic flag to false, encouraging discovery of new optimal actions. During normal operation, the system defaults to exploitative mode, prioritizing actions with known high rewards [20].
@@
This context-sensitive strategy ensures that the system balances learning new responses during anomalous conditions with reliable execution during normal operations, avoiding both excessive exploration that could disrupt stable processes and insufficient exploration that would prevent adaptation to novel situations.
@@ III-D3 Dynamic Rule Updating
The Adaptation Layer continuously refines the symbolic knowledge base by incorporating new rules derived from reinforcement learning outcomes. This dynamic updating process ensures that the symbolic reasoning component evolves alongside changing operational conditions and user preferences, maintaining alignment between the system’s explicit knowledge representation and its learned policies. The integration of RL-derived insights into symbolic rules creates a feedback loop where improved policies inform better rules, which in turn guide more effective policy learning.
@@ III-E Framework Integration and Adaptive Operation
ANSR-DT implements tight integration across its three layers to enable seamless information flow and adaptive decision-making. The operational sequence illustrated in Fig. 2 shows how sensor data progresses through preprocessing, neural pattern extraction, symbolic reasoning, and adaptive control generation in a continuous cycle. This integration maintains bidirectional communication between physical and digital components through a data preprocessing pipeline and feedback control loop.
The data preprocessing pipeline systematically validates and normalizes sensor data before feeding into the processing layer. Sensor readings undergo standardization using rolling statistics to ensure consistent scaling across different measurement types. Multi-sensor data streams are temporally aligned to account for varying sampling rates across temperature, vibration, and pressure sensors. The pipeline identifies deviations in raw sensor readings that may indicate measurement faults or genuine operational anomalies requiring investigation. This preprocessing ensures that the neural component receives clean, aligned inputs suitable for temporal pattern extraction.
The feedback control loop provides real-time optimization of the sensor network itself. The processing layer dynamically adjusts sampling rates based on current system state, increasing measurement frequency when potential anomalies are detected and reducing it during stable operation to conserve resources. Automated sensor calibration occurs based on performance metrics derived from prediction accuracy and consistency checks across redundant measurements. The system continuously performs fault detection and sensor reliability assessment, flagging degraded sensors for maintenance before they compromise monitoring effectiveness.
Continuous adaptation operates through an integrated feedback loop encompassing state assessment and adaptive response. The system continuously monitors performance metrics including accuracy, latency, and resource utilization. Efficiency indices track how effectively the system uses computational and sensing resources. Rule activation patterns and effectiveness are analyzed to identify which symbolic rules provide the most value in different operational contexts. Policy performance is evaluated across operational modes to ensure that learned behaviors generalize appropriately. Based on this assessment, the system coordinates adaptive responses including policy refinement through PPO-based learning, knowledge base evolution via dynamic rule extraction, and sensor network configuration optimization. This integration ensures that ANSR-DT maintains optimal performance while adapting to changing operational conditions, with each layer contributing to the system’s overall adaptability and robustness [28].